An office has 30 computers. Seventeen of the 30 are Macintosh, and the remaining thirteen are windows. Two computers are randomly selected without replacement.

What is the probability of randomly selecting a Macintosh given that the first computer selected was a windows machine? If needed, round to FOUR decimal places.

Pr(Macintosh|Windows) = ___________
Answer #1

Macintosh = 17

Windows = 13

Total = 30

Selection 1: WINDOWS

P(Selection 1: WINDOWS) = 13/30 = 0.4333

Selection 2: Macintosh

Macintosh = 17

Windows= 12

Total = 29

P(Selection 2: Macintosh) = 17/29 = 0.5862

So,

Answer is:

0.5862